OpenStudy (anonymous):

A six-member committee is formed to plan a Christmas formal. The members of the group are to be selected from five teachers and nine students. If the members are randomly selected, what is the probability that it will include at least two teachers?

OpenStudy (anonymous):

Teacher = T, Student = S (5T,1S) + (4T,2S) + (3T,3S) + (2T,4S) now find probability of each and add them up

OpenStudy (anonymous):

That is the part i knew how to do lol what is the probability equation?

OpenStudy (anonymous):

P(at least 2 teachers)=1-P(0 teachers)-P(1 teacher) There are 14C6 ways to a committee of 6 from 14. P(no teachers)=(9C6)(5C0)/(14C6) P(1 teacher)=(9C5)(5C1)/(14C6) So P(at least 2 teachers)=1-(9C6)(5C0)/(14C6)-(9C5)(5C1)/(14C6) =2289/3003=about 0.7622 or 76.22%